OpenStudy (mendicant_bias):

Related Rates challenge problem, I think I understand it but I can't help but feel that the answer itself is slightly wrong. (One minute.)

OpenStudy (mendicant_bias):

This is what the answer is supposed to be, which I think may be wrong:

OpenStudy (anonymous):

why? i thind the develoment of the problem is right

OpenStudy (mendicant_bias):

This is what I did: \[A = xy\] \[\frac{ dA }{ dt } = y (\frac{ dy }{ dt }) + x (\frac{ dx }{ dt })\] \[f(x) = x ^{3}-2xy+y ^{3}+1 = 0\] \[f'(x) = 3x ^{2}(\frac{ dx }{ dt }) -2y (\frac{ dx }{ dt }) - 2x (\frac{ dy }{ dt })+3y ^{2}(\frac{ dy }{ dt }) = 0\] One sec, coming back to this.

OpenStudy (mendicant_bias):

I think the development is correct, too, it's just at the very end that I take issue. Shouldn't, at the end, since we're ultimately looking for dA/dt, we plug back into that? After finding dy/dt, and given the x, y, and dx/dt values, shouldn't we plug back in? They find the value of dy/dt and then say that that is dA/dt.

OpenStudy (anonymous):

yes there is a step that they r missing

OpenStudy (mendicant_bias):

Made a mistake in rewriting out the last derivative, but that doesn't change my point. But yeah, it can't be that, right? It's hopefully just a writing error.

OpenStudy (mendicant_bias):

\[f'(x) = 3x ^{2}(\frac{ dx }{ dt }) -2y ^{2}(\frac{ dx }{ dt })-4xy(\frac{ dy }{ dt })+3y ^{2}(\frac{ dy }{ dt }) = 0\] \[\frac{ dy }{ dt }=\frac{ (3x ^{2}-2y ^{2})(\frac{ dx }{ dt }) }{ 4xy-3y ^{2} }\] Plugging in now for dy/dt = 1, x = 2, y = 3: \[\frac{ dy }{ dt }=\frac{ [3(2)^{2}-2(3)^{2}](1) }{ 4(2)(3)-3(3)^{2} }\] \[\frac{ dy }{ dt }=\frac{ -6 }{ -3 } = 2\] Whoah, so, their dy/dt value is wrong, too? Jesus.

OpenStudy (mendicant_bias):

\[\frac{ dA }{ dt }= (1)(3)+(2)(2)=7\]Wait, so....the dA/dt value they specified is correct...it's just the dy/dt value that they fudged. Am I seeing this right?

OpenStudy (phi):

looks ok except you have a typo and want dy/dt = 2 and you should put in one more line to show dA/dt = 7

OpenStudy (amistre64):

Area of rectangle\[A=xy\] \[A'=x'y+xy'\]they give you x', x, and y \[A'=1(3)+2y'\] use the controlling function (it controls the x and y values) to find y' \[x^3-2xy^2+y^3+1=0\] \[3x^2x'-2x'y^2-4xyy'+3y^2y'+0=0\] x'=1, x=2, y=3 \[3(2)^2-2(3)^2-4(2)(3)y'+3(3)^2y'+0=0\] solve for y' \[12-18-24y'+27y'=0\] \[12-18+3y'=0\] \[4-6+y'=0\] y' = 2 plugging that into the Area we get:\[A'=1(3)+2(2)=7\]

OpenStudy (mendicant_bias):

@phi: I did, I just put it in the next post, and I did write that dy/dt = 2.Thanks, though. guys.